FOURIER TRANSFORM ELECTRONIC SPECTROSCOPY OF STRONTIUM HYDRIDE: SrH and SrD

\underline{KALEY A. WALKER}, \underline{TREANA PAREKUNNEL}, KEITH TERESZCHUK, IOULI GORDON AND PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1.

During a search for double minima states of several alkaline earth hydrides, high resolution spectra of the E^{2}Π − X^{2}Σ^{+} transitions of these species were recorded. The results for SrH and SrD, in the 18500 and 19500 cm^{−1} region, are presented in this paper. A new combined discharge and tube furnace source was used to generate the molecules. Strontium metal was heated in an alumina tube to 650^{∘}C with a slow flow of Ar (∼ 20 torr) and H_{2}/D_{2} (∼ 1 torr), and a 3kV/330 mA continuous DC discharge was applied. The spectra were recorded using a Fourier transform spectrometer. The resulting molecular constants for SrH and SrD will be presented and discussed.